OpenStudy (anonymous):

The length of a rectangle is 8 meters more than its width. If the area is 65 square meters, what are the length and width?

OpenStudy (anonymous):

Let x= width Let y = length y = 8 + x Area = length x width A = (8+x)(x) 65 = (8+x)(x) 65 = 8x+x^2 0 = x^2 + 8x -65 I don't think it works out perfectly but try solving that

OpenStudy (anonymous):

hmm...

OpenStudy (anonymous):

would it be (x+13)(x-5) or no?

OpenStudy (anonymous):

Yes! Haha I didnt think it'd be a perfect polynomial but it does work out! The width is 5 (since the width cant be negative) and then the length is 13.